# Gravitational Atoms from Topological Stars

### v1

@@ III.2 Zero-KK modes
Figs. (5) and (6) showcase the validity of our analysis for two different representative sets of parameters in the case of a TS of the second and first kind, respectively. In both cases we see that the $\ell=0$, $n=0$ mode is the one with the highest binding energy, $\gamma$ (that is - lowest frequency $\tilde{\omega}$). As we increase $\ell$, but keep the radial overtone number fixed, the energy decreases (the frequency increases) and from the eikonal approximation, (C), we know that it approaches zero asymptotically (the frequency tends to the scalar mass, $\tilde{\mu}$, in the same limit, for $\tilde{k}=0$). Moreover, the $n=0$ modes almost saturate the bounds on their binding energies in all three regions ($\gamma_{0}$ for TS modes, and $\gamma_{+}$ for mixed and cloud modes). As the overtone number, $n$, is increased, while keeping $\ell$ fixed, the energies decrease, faster for the TS modes in comparison to the cloud ones (for mixed modes the rate depends on which type dominates the "mixture"). We expect that for large overtones, $n\gg 1$, the binding energy will tend to zero, $\gamma\to 0$. This picture holds qualitatively for all scalar masses, $\tilde{\mu}$, and $\eta$ that we have probed, except for $\sigma\to 0$ (for $\tilde{k}=0$ this is equivalent to $\tilde{\mu}\ll 1$), when the TS region shrinks to an infinitesimal size.
@@
Finally, in the plots, we have also included the binding energy associated to the circular timelike geodesics from Section (II.2): $\gamma_{g}=\sqrt{\tilde{\mu}^{2}\,\big(1-\mathcal{E}^{2}\big)}$ with $\mathcal{E}$ given in (36). The analysis in Section (II.2) holds to leading order in the eikonal limit, and, as expected, $\gamma_{g}$ agrees with $\gamma_{+}$ to that order for $\ell\gg 1$.
@@
For the $\eta=1.75$ case in Fig. 8, the picture is essentially the same, albeit with some slight qualitative differences in the $\ell=7$ mixed-mode case due to the second potential minimum being deeper than the potential’s value at the star’s surface.
Figure 7: Top Panels: Scalar wavefunctions on the topological star background with parameters as in Fig. 5, with increasing overtone number. Bottom Panels: $V_{\rm eff}$ (Black line) for the chosen parameters compared with overtone “energies” $\epsilon$ (colored lines). We take $\ell=8$, $\ell=10$, $\ell=12$, $\ell=15$, corresponding to the $\ell<\Lambda_{-}$, $\Lambda_{-}<\ell<\Lambda_{+}$, $\Lambda_{+}<\ell<\Lambda_{0}$, and $\Lambda_{0}<\ell$ regions of parameter space, respectively. The gray line corresponds to the ISCO radius $3r_{S}$. Modes with low $\ell$ always localize close to the star. Waves localize further from the star’s surface as both overtone number and $\ell$ values increase, in line with the mode classification scheme in Eqs. (LABEL:eq:TSregion)–(LABEL:eq:cloudRegion).
Figure 8: Same as Fig. 7, except taking the parameters in Fig. 6. From left to right, we take $\ell=5$, $\ell=7$, $\ell=9$, corresponding to the $\ell<\Lambda_{-}$, $\Lambda_{0}<\ell<\Lambda_{+}$, and $\Lambda_{+}<\ell$ regions, respectively. Note that there is no $\ell$ value which lies within the $\Lambda_{-}<\ell<\Lambda_{0}$ region for the parameters chosen in Fig. 6.
The transitions from TS modes to mixed-modes and cloud modes as angular momentum increases is further demonstrated in Fig. 9. We show how the fundamental mode wavefunctions change as $\ell$ increases near the threshold region $\ell\sim\Lambda_{-}$ for the same parameters as Figs. 5 and 6. In both cases, we see the waves localize near the star, and gradually transition to localizing in the cloud region. Physically, this may be explained by noting that larger angular momentum in the field serves to push the field further outwards, towards the region where the field exhibits massive geodesic behavior.
Figure 9: Left: Wavefunction and potential for the fundamental mode, taking Fig. 5’s parameters and varying $\ell$. Near the transition region at $\ell\sim 9$, the wavefunctions begin to move outwards towards the cloud, and eventually localize in the cloud once the $\ell>\Lambda_{+}$ threshold is reached. Right: Same as the left panel, for Fig. 6’s parameters. In both figures, the location of the ISCO radius is represented by a gray line.
As advanced in the analysis of Section II.2, in the limit $\mu r_{b}\gg 1$ the spectrum is determined by timelike geodesics. There, it was shown that for large enough $\ell$, in particular $\ell\gtrsim\sqrt{3}\mu r_{s}$, the mode frequencies are dictated by the energy of stable circular geodesics,
@@
where the function $\mathcal{E}(x)$ is given in (36). This family terminates at $\ell\sim\sqrt{3}\mu r_{s}$, below which the solutions are no longer localized around circular geodesics, and, modes probe the interior of the ISCO. This behavior is illustrated in Fig. 10, where the geodesic prediction (64) is contrasted with the accurate numerical results for $\mu r_{b}=50$, which we expect to be well within the regime $\mu r_{b}\gg 1$. One can see that the curve given by (64) follows very well the numerical result, and, terminates at $\ell=\sqrt{3}\mu r_{s}$, which is with good precision the angular momentum where the solutions change behavior from cloud to TS modes.
Figure 10: Fundamental normal mode spectrum in the regime $\mu r_{b}\gg 1$. It can be seen that the circular geodesic prediction (64), represented in black by $\gamma_{g}=\sqrt{\tilde{\mu}^{2}\,\big(1-\mathcal{E}^{2}\big)}$ with $\mathcal{E}$ from (36), describes well the behavior of cloud modes, and terminates with good precision where cloud modes transition into TS ones.
Finally, in Fig. 11 we compare the hydrogenic ($\tilde{\mu}\ll 1$) limit prediction in Eq. (3) to numerical results obtained through Leaver’s method, setting $k=\ell=0$. We find good agreement with the prediction as $\tilde{\mu}$ is made smaller, and find qualitatively similar results for other choices of $k,\ell,$ and $r_{b}$.
Figure 11: Comparison of normal mode spectra for various choices of $\tilde{\mu}$, in the hydrogenic limit $\tilde{\mu}\ll 1$. Solid lines are the analytic prediction Eq.(3), and points are the numerical results generated through Leaver’s method. We find good agreement with the predictions of Eq. (3), which improves as the value of $\tilde{\mu}$ is made smaller. Here we set $\eta=1.25$, $\ell=\tilde{k}=0$ to demonstrate the spectrum’s behavior, but find qualitatively similar results for nonzero choices of KK momentum, and other values of $\eta,\ell$.
@@ III.3 KK modes
Allowing the scalar field to carry KK momentum does not lead to any major qualitative differences with the picture from the previous subsection. However, looking at the definition of Q(r~)Q($\tilde${r}) in (III), we see that performing a similar analysis is obstructed by the increase in degree in $\tilde{r}$ of the polynomial we have to work with. It turns out that it is slightly easier to include the prefactors and consider:
@@
One can then technically replicate what was done for $\tilde{k}=0$, however, we also need to look at the second and third derivatives of (65) and almost all steps of the process are more complicated with more subcases to consider, resulting in a lot of cumbersome expressions. We will therefore only mention two conditions that nonetheless allow one to, at least, distinguish the pure cloud modes from the rest.
@@
[…]
@@
It is possible for $\gamma_{\tilde{k}}>\sigma_{\eta}/3$ for $\Lambda_{\tilde{k}}<\sigma_{\eta}$. When that happens, that part of the cloud region is not present.
There are 2 differences with the $\tilde{k}=0$ scenario. First,
@@
leading to $\gamma_{\tilde{k}}$ being complex and, thus, the TS region disappearing for any $\ell$. This is in contrast to the $\tilde{k}=0$ case, where, at least, for $\ell=0$ we are always guaranteed some non-zero area for the TS region, as discussed shortly after (62). Taking the definition of $\sigma$, (43), and the requirement that it is positive for the existence of bound modes, allows us to rewrite (70) in terms of the scalar mass:
@@
Figure 12: Binding energy, $\gamma$, (49), of massive, scalar field modes as a function of $\ell$, in the TS case, for $\eta=1.75$, $\tilde{\mu}=7$, $N=0$, evaluated numerically for the n=0, 1, 2, 3n=0,\,1,\,2,\,3 radial overtones in cyan disks, orange diamonds, pink pentagons and green six-pointed stars, respectively, with $\tilde{k}=1,\,9,\,12,\,15\times\frac{1}{2}\sqrt{\frac{\eta-1}{\eta}}$ KK momentum carried by the scalar for the top left, top right, bottom left and bottom right plots, respectively. Log plots for clarity. Filled markers are pure TS modes. Empty markers correspond to cloud modes, localized outside of the timelike geodesics ISCO. Half-filled markers indicate mixed modes. The respective regions, where our analysis specifies they can exist are shaded differently for clarity. The dashed green line for the eikonal approximation, (C), and the dashed-dotted magenta line for the large scalar mass approximation, (D), are for $n=0$ only.
The latter novelty is demonstrated in Fig. (12), where we have shown the equivalent of Fig. (6), but for non-zero value of KK momentum carried by the scalar. We remind the reader that the latter is quantized, $k=\frac{p}{R_{y}}$, where $p\in\mathbb{N}$ and $R_{y}$ is the radius of the extra compact dimension, subject to the regularity condition (9). In our dimensionless [^8] We have also set $R_{y}=1$. units:
@@
Moreover, we have only looked at the case of orbifold parameter $K=1$ in this work, hence we label the KK momentum as given by $p$ units of $\frac{1}{2}\sqrt{\frac{\eta-1}{\eta}}$, and we only indicate the value of $p$ on the plots. For $\eta=1.75$ and $\tilde{\mu}=7$, (71) gives us a maximum of $p=24$ units allowed and a transition to a situation with no TS modes for at least $p=15$ units. Indeed, on the bottom right graph of Fig. (12), where the case of $p=15$ is shown, there is no TS region and we have only found cloud modes. Moreover, at least for these parameters, the mixed region shrinks very quickly. It is barely visible on the top left plot in Fig. (12), which is otherwise not much different than Fig. (6), and demonstrates the effect of the minimum amount of KK momentum allowed. For $p=9$ and $p=12$ there is no mixed region at all. Of course, once the TS region is gone, so is the mixed one. We do not have an analytical expression that we can refer to in order to determine the conditions for its shrinking. The regions in Fig. (12) were determined numerically.
In addition, one can also see in Fig. (12) how the large scalar mass approximation, (D), loses validity with increasing $\tilde{k}$. This is expected, as (D), assumes that all parameters are $\mathcal{O}(1)$ with respect to $\tilde{\mu}\gg 1$, and the plots on the bottom row already have $\tilde{k}\sim\tilde{\mu}$. Furthermore, the approximation, (D), is designed to capture TS modes - localized near the star surface, thus, once the TS region is gone, it should break down.
@@
Next, we present the equivalent of Fig. (7), but this time we fix $\ell=10$ and vary the amount of scalar field’s KK momentum on the different plots in Fig. (13). This serves to show how the increase in $\tilde{k}$ slowly pushes all the modes out, until we have only cloud ones left. For $\eta=1.25$ and $\tilde{\mu}=7$, the parameters in Fig. (7), the bounds in (71) are: maximum of $p=62$ and no TS region for $p=37$ and higher. For small values of KK momentum, as on the leftmost figure, the qualitative features of the $\tilde{k}=0$ case are still visible: as we increase the number of overtones we transition from TS to mixed and back to TS modes. The first transition does happen earlier - at $n=2$ in comparison to $n=4$ in Fig. (7). Further increasing $\tilde{k}$, middle plot, turns the initial TS modes into cloud ones (due to the TS region shrinking) with the rest of the transitions as before, until finally on the rightmost graph there are no more TS or mixed modes.
@@ III.4 Charged modes and gravitating Thomson dipoles
In this section we discuss the case in which the scalar field is electrically charged. We remind that, for regularity reasons, the field’s fundamental charge $e$ and the star’s magnetic charge $P$ are related by Dirac’s quantization condition $eP=N/2$, with $N$ any integer (see (11)). We chose to work in terms of $P$ and $N$ as independent variables (thus fixing $e=N/2P$), so that $N=0$ corresponds to the neutral case, $N=\pm 1$ to the 1-monopole star, $N=\pm 2$ to the 2-monopole star, etc. The electromagnetic coupling enters the radial equation (15) only through the constant $\Lambda$, given by
@@
so the mode classification and other analysis about the radial function and the normal frequencies of Sections III.2 and III.3 also accounts for charged scalars. In certain aspects, charged and neutral modes are almost identical. For example, two normal modes with the same $\tilde{\mu}$ and $\Lambda$ will possess identical radial profiles (up to global normalization) and frequency spectrum, regardless of the values that $N$ and respective $\ell$’s may take. However, regarding other characteristics they can be strikingly different. Even if a neutral and a charged mode have identical frequencies, radial profiles and total energies, they exhibit distinct spatial distributions due to the difference between Wu-Yang monopole harmonics and ordinary spherical harmonics (see Appendix A). Consequently, such two modes possess different amounts of angular momentum, with the charged field consistently capable of carrying a larger value.
@@
where we set $R_{y}=1$ and restricted to zero-KK modes for simplicity. Consider now two normal modes with equal masses $\tilde{\mu}$, but one being neutral $\varphi_{\text{n}}$ (so $N_{\text{n}}=0$) and the other charged $\varphi_{\text{c}}$ (so $N_{\text{c}}\neq 0$) and possessing different angular numbers $\ell_{\text{n}}$ and $\ell_{\text{c}}$, respectively. Furthermore, assume that the charge number $N_{\text{c}}$ can be taken such that $\Lambda_{\text{n}}=\Lambda_{\text{c}}$, where in general $\Lambda$ is given by (73) (such $N_{\text{c}}$ does not always exist for arbitrary values of $\ell_{\text{n,c}}$, but it is convenient to restrict to this case in this discussion). Two such modes have the same frequency spectrum, and their radial wavefunctions are also identical if we rescale them suitably to impose that they have the same total energy, $E[\varphi_{\text{n}}]=E[\varphi_{\text{c}}]$. However, recall that the angular numbers must satisfy
@@
From the fact that $\Lambda_{\text{n}}=\Lambda_{\text{c}}$ it is clear that $\ell_{\text{n}}<\ell_{\text{c}}$, so in the charged case $\lvert m_{\text{c}}\rvert$ can take larger values than $\lvert m_{\text{n}}\rvert$. Thus, the charged field can carry more angular momentum as follows from (75). In particular, focusing on the field states that possess a maximum amount of angular momentum, $\ell_{\text{n,c}}=m_{\text{n,c}}$, we conclude that
@@
To explain this difference, it is useful to illustrate how $\varphi_{\text{n,c}}$ distribute in space. For concreteness, we shall restrict ourselves to fundamental modes with zero KK momentum and take the same background parameters and field mass of Fig.6 ($\eta=1.75$, $\tilde{\mu}=7$). For $\varphi_{\text{n}}$, we shall take maximum angular momentum states with $\ell_{\text{n}}=m_{\text{n}}=6,7,8$, which according to the mode classification given in Section III.2 correspond to TS, mixed and could modes, respectively. To guarantee that $\Lambda_{\text{n}}=\Lambda_{\text{c}}$, in the charged case we also take maximum angular momentum states with $\ell_{\text{c}}=m_{\text{c}}=N_{\text{c}}/2$ and $N_{\text{c}}=84,112,144$. We emphasize that the condition $m_{\text{c}}=N_{\text{c}}/2$ corresponds to a north monopole mode, that is, a mode that localizes along the north component of the $z$-axis and carries angular momentum in the $z$-direction Pereñiguez et al. (2024) (see also Appendix A). In all cases, we normalize the wavefunctions by setting the total energy to $E[\varphi_{\text{n,c}}]/r_{B}=10^{-3}$.

### v2

@@ III.2 Zero-KK modes
In Figs. 5 and 6 we showcase the validity of our analysis for two different representative sets of parameters in the case of a TS of the second and first kind, respectively. In both cases we see that the $\ell=0$, $n=0$ mode is the one with the highest binding energy $\gamma$ (i.e., the lowest frequency $\tilde{\omega}$). As we increase $\ell$, but keep the radial overtone number fixed, the energy decreases (the frequency increases) and from the eikonal approximation of Eq. (C), we know that it approaches zero asymptotically (the frequency tends to the scalar mass $\tilde{\mu}$ in the same limit, for $\tilde{k}=0$). Moreover, the $n=0$ modes almost saturate the bounds on their binding energies in all three regions ($\gamma_{0}$ for TS modes, and $\gamma_{+}$ for mixed and cloud modes). As the overtone number $n$ is increased, while keeping $\ell$ fixed, the energies decrease, faster for the TS modes in comparison to the cloud ones (for mixed modes the rate depends on which type dominates the “mixture”). We expect that for large overtones, $n\gg 1$, the binding energy will tend to zero, $\gamma\to 0$. This picture holds qualitatively for all scalar masses $\tilde{\mu}$ and for all $\eta$’s that we have probed, except for $\sigma\to 0$ (for $\tilde{k}=0$ this is equivalent to $\tilde{\mu}\ll 1$), when the TS region shrinks to an infinitesimal size.
@@
Finally, in the plots, we have also included the binding energy associated to the circular timelike geodesics from Section II.2: $\gamma_{g}=\sqrt{\tilde{\mu}^{2}\,\big(1-\mathcal{E}^{2}\big)}$, with $\mathcal{E}$ given in Eq. (36). The analysis in Section II.2 holds to leading order in the eikonal limit, and, as expected, $\gamma_{g}$ agrees with $\gamma_{+}$ to that order for $\ell\gg 1$.
@@
For the $\eta=1.75$ case in Fig. 8, the picture is essentially the same, albeit with some slight qualitative differences in the $\ell=7$ mixed-mode case, due to the second potential minimum being deeper than the potential’s value at the star’s surface.
Figure 7: Top Panels: Scalar wavefunctions on the TS background with parameters as in Fig. 5, with increasing overtone number. Bottom Panels: $V_{\rm eff}$ (Black line) for the chosen parameters compared with overtone “energies” $\epsilon$ (colored lines). We take $\ell=8$, $\ell=10$, $\ell=12$, $\ell=15$, corresponding to the $\ell<\Lambda_{-}$, $\Lambda_{-}<\ell<\Lambda_{+}$, $\Lambda_{+}<\ell<\Lambda_{0}$, and $\Lambda_{0}<\ell$ regions of parameter space, respectively. The gray line corresponds to the ISCO radius $3r_{S}$. Modes with low $\ell$ are always localized close to the star. Waves are localized further from the star’s surface as both overtone number and $\ell$ values increase, in line with the mode classification scheme in Eqs. (LABEL:eq:TSregion)–(LABEL:eq:cloudRegion).
Figure 8: Same as Fig. 7, but for the parameters in Fig. 6. From left to right, we take $\ell=5$, $\ell=7$, $\ell=9$, corresponding to the $\ell<\Lambda_{-}$, $\Lambda_{0}<\ell<\Lambda_{+}$, and $\Lambda_{+}<\ell$ regions, respectively. Note that there is no $\ell$ value which lies within the $\Lambda_{-}<\ell<\Lambda_{0}$ region for the parameters chosen in Fig. 6.
The transitions from TS modes to mixed modes and cloud modes as the angular momentum increases is further demonstrated in Fig. 9. We show how the fundamental mode wavefunctions change as $\ell$ increases near the threshold region $\ell\sim\Lambda_{-}$ for the same parameters as in Figs. 5 and 6. In both cases, we see the waves localize near the star, and gradually transition to localizing in the cloud region. Physically, this may be explained by noting that larger angular momentum in the field serves to push the field further outwards, towards the region where the field exhibits massive geodesic behavior.
Figure 9: Left: Wavefunction and potential for the fundamental mode, using the same parameters as in Fig. 5 and varying $\ell$. Near the transition region at $\ell\sim 9$, the wavefunctions begin to move outwards towards the cloud, and eventually localize in the cloud once the $\ell>\Lambda_{+}$ threshold is reached. Right: Same as the left panel, for for the parameters used in Fig. 6. In both figures, the location of the ISCO radius is represented by a gray vertical line.
As anticipated in the analysis of Section II.2, in the limit $\mu r_{b}\gg 1$ the spectrum is determined by timelike geodesics. There, it was shown that for large enough $\ell$, in particular $\ell\gtrsim\sqrt{3}\mu r_{s}$, the mode frequencies are dictated by the energy of stable circular geodesics
@@
where the function $\mathcal{E}(x)$ is given in Eq. (36). This family terminates at $\ell\sim\sqrt{3}\mu r_{s}$, below which the solutions are no longer localized around circular geodesics, and modes probe the interior of the ISCO. This behavior is illustrated in Fig. 10, where the geodesic prediction (64) is contrasted with accurate numerical results for $\mu r_{b}=50$, which we expect to be well within the regime $\mu r_{b}\gg 1$. One can see that the curve given by (64) follows very well the numerical result and terminates at $\ell=\sqrt{3}\mu r_{s}$, which is with good precision the angular momentum where the solutions change behavior from cloud to TS modes.
Figure 10: Fundamental normal mode spectrum in the regime $\mu r_{b}\gg 1$. The circular geodesic prediction (64), represented in black by $\gamma_{g}=\sqrt{\tilde{\mu}^{2}\,\big(1-\mathcal{E}^{2}\big)}$ with Eq. $\mathcal{E}$ from (36), describes well the behavior of cloud modes, and terminates with good precision where cloud modes transition into TS modes.
Finally, in Fig. 11 we compare the hydrogenic ($\tilde{\mu}\ll 1$) limit prediction in Eq. (3) to numerical results obtained through Leaver’s method, setting $k=\ell=0$. We find good agreement with the prediction as $\tilde{\mu}$ decreases. We find qualitatively similar results for other choices of $k,\ell,$ and $r_{b}$.
Figure 11: Comparison of normal mode spectra for various choices of $\tilde{\mu}$, in the hydrogenic limit $\tilde{\mu}\ll 1$. Solid lines are the analytic prediction of Eq. (3), and points are the numerical results generated through Leaver’s method. We find good agreement with the predictions of Eq. (3), which improves as the value of $\tilde{\mu}$ is made smaller. Here we set $\eta=1.25$, $\ell=\tilde{k}=0$ to demonstrate the behavior of the spectrum, but we find qualitatively similar results for nonzero choices of KK momentum and for other values of $\eta$ and $\ell$.
@@ III.3 KK modes
Allowing the scalar field to carry KK momentum does not lead to any major qualitative changes from the picture outlined in the previous subsection. However, looking at the definition of Q(r~)Q($\tilde${r}) in (III), we see that performing a similar analysis is obstructed by the increase in degree in $\tilde{r}$ of the polynomial we have to work with. It turns out that it is slightly easier to include the prefactors and consider:
@@
One can then technically replicate what was done for $\tilde{k}=0$, however, we also need to look at the second and third derivatives of Eq. (65) and almost all steps of the process are more complicated with more subcases to consider, resulting in a lot of cumbersome expressions. We will therefore only mention two conditions that nonetheless allow one to, at least, distinguish the pure cloud modes from the rest.
@@
$\varphi_{\tilde{k}}$ is given by Eq. (43), implying $\gamma_{\tilde{k}=0}=\gamma_{0}$, $\Lambda_{\tilde{k}}\geq 0$ for any choice of parameters, and $\Lambda_{\tilde{k}=0}=\Lambda$. The reality of the square root inside $\gamma_{c}$ requires $\Lambda_{\tilde{k}}>\sigma_{\eta}/2$, which is always satisfied in the regime of interest [^5] Note that $\gamma_{c}$ does not reduce to $\gamma_{+}$ for $\tilde{k}=0$, because we derive the two bounds in different ways. In particular, we expect that the equivalent of $\gamma_{+}$ for non-zero $\tilde{k}$ will provide a more stringent bound than $\gamma_{c}$.. The two regions thus defined can overlap and the area where that happens can host any type of mode – pure or mixed. Further analysis of $Q_{k}$ and its derivatives can allow one to precisely characterize how that occurs, but the conditions are hard to work with. On the other hand, we can easily define the regions where only one set of modes exists:
@@
It is possible for $\gamma_{\tilde{k}}>\sigma_{\eta}/3$ when $\Lambda_{\tilde{k}}<\sigma_{\eta}$. When that happens, that part of the cloud region is not present.
There are two differences with the $\tilde{k}=0$ scenario. First,
@@
leading to $\gamma_{\tilde{k}}$ being complex and, thus, the TS region disappearing for any $\ell$. This is in contrast to the $\tilde{k}=0$ case, where at least for $\ell=0$ we are always guaranteed some non-zero area for the TS region, as discussed below Eq. (62). Taking the definition of $\sigma$, Eq. (43), and the requirement that it is positive for the existence of bound modes, allows us to rewrite Eq. (70) in terms of the scalar mass:
@@
Figure 12: Same as Fig. 5, but for $\eta=1.75$, $\tilde{\mu}=7$, $N=0$. Now the scalar carries KK momentum $\tilde{k}=1,\,9,\,12,\,15\times\frac{1}{2}\sqrt{\frac{\eta-1}{\eta}}$ (top left, top right, bottom left and bottom right plots, respectively).
This novelty is demonstrated in Fig. 12, where we have shown the equivalent of Fig. 6, but for non-zero value of KK momentum carried by the scalar. We remind the reader that the latter is quantized, $k=\frac{p}{R_{y}}$, where $p\in\mathbb{N}$ and $R_{y}$ is the radius of the extra compact dimension, subject to the regularity condition (9). In our dimensionless units (and also setting $R_{y}=1$):
@@
Moreover, we have only looked at the case of orbifold parameter $K=1$ in this work, hence we label the KK momentum as given by $p$ units of $\frac{1}{2}\sqrt{\frac{\eta-1}{\eta}}$, and we only indicate the value of $p$ on the plots. For $\eta=1.75$ and $\tilde{\mu}=7$, Eq. (71) gives us a maximum of $p=24$ units allowed and a transition to a situation with no TS modes for at least $p=15$ units. Indeed, on the bottom right graph of Fig. 12, where the case of $p=15$ is shown, there is no TS region and we have only found cloud modes. Moreover, at least for these parameters, the mixed region shrinks very quickly. It is barely visible on the top left plot in Fig. 12, which is otherwise not much different than Fig. 6, and demonstrates the effect of the minimum amount of KK momentum allowed. For $p=9$ and $p=12$ there is no mixed region at all. Of course, once the TS region is gone, so is the mixed one. We do not have an analytical expression that we can refer to in order to determine the conditions for its shrinking. The regions in Fig. 12 were determined numerically.
In addition, one can also see in Fig. 12 how the large scalar mass approximation of Eq. (D) loses validity with increasing $\tilde{k}$. This is expected, as Eq. (D) assumes that all parameters are $\mathcal{O}(1)$ with respect to $\tilde{\mu}\gg 1$, and the plots on the bottom row already have $\tilde{k}\sim\tilde{\mu}$. Furthermore, the approximation (D) is designed to capture TS modes, which are localized near the star surface, and thus it should break down once the TS region is gone.
@@
Next, we present the equivalent of Fig. 7, but this time we fix $\ell=10$ and vary the amount of the scalar field’s KK momentum on the different plots in Fig. 13. This serves to show how the increase in $\tilde{k}$ slowly pushes all the modes out, until we have only cloud ones left. For $\eta=1.25$ and $\tilde{\mu}=7$, the parameters in Fig. 7, the bounds in Eq. (71) are: maximum of $p=62$ and no TS region for $p=37$ and higher. For small values of KK momentum, as on the leftmost figure, the qualitative features of the $\tilde{k}=0$ case are still visible: as we increase the number of overtones we transition from TS to mixed and back to TS modes. The first transition does happen earlier, at $n=2$ in comparison to $n=4$ in Fig. 7. Further increasing $\tilde{k}$ (middle plot) turns the initial TS modes into cloud ones (due to the TS region shrinking) with the rest of the transitions as before, until finally on the rightmost plot there are no more TS or mixed modes.
@@ III.4 Charged modes and gravitating Thomson dipoles
In this section we discuss the case in which the scalar field is electrically charged. We remind the reader that, for regularity reasons, the field’s fundamental charge $e$ and the star’s magnetic charge $P$ are related by Dirac’s quantization condition $eP=N/2$, with $N$ any integer: see Eq. (11). We choose to work in terms of $P$ and $N$ as independent variables (thus fixing $e=N/2P$), so that $N=0$ corresponds to the neutral case, $N=\pm 1$ to the 1-monopole star, $N=\pm 2$ to the 2-monopole star, etc. The electromagnetic coupling enters the radial equation (15) only through the constant $\Lambda$, given by
@@
so the mode classification and other analysis about the radial function and the normal frequencies of Sections III.2 and III.3 also accounts for charged scalars. In certain aspects, charged and neutral modes are almost identical. For example, two normal modes with the same $\tilde{\mu}$ and $\Lambda$ will possess identical radial profiles (up to global normalization) and frequency spectrum, regardless of the values that $N$ and the respective $\ell$’s may take. However, regarding other characteristics they can be strikingly different. Even if a neutral and a charged mode have identical frequencies, radial profiles and total energies, they exhibit distinct spatial distributions due to the difference between Wu-Yang monopole harmonics and ordinary spherical harmonics (see Appendix A). Consequently, such two modes possess different amounts of angular momentum, with the charged field consistently capable of carrying a larger value.
@@
where we have set $R_{y}=1$ and restricted to zero-KK modes for simplicity. Consider now two normal modes with equal masses $\tilde{\mu}$, but one being neutral $\varphi_{\text{n}}$ (so $N_{\text{n}}=0$) and the other charged $\varphi_{\text{c}}$ (so $N_{\text{c}}\neq 0$) and possessing different angular numbers $\ell_{\text{n}}$ and $\ell_{\text{c}}$, respectively. Furthermore, assume that the charge number $N_{\text{c}}$ can be taken such that $\Lambda_{\text{n}}=\Lambda_{\text{c}}$, where in general $\Lambda$ is given by (73) (such $N_{\text{c}}$ does not always exist for arbitrary values of $\ell_{\text{n,c}}$, but it is convenient to restrict to this case in this discussion). Two such modes have the same frequency spectrum, and their radial wavefunctions are also identical if we rescale them suitably to impose that they have the same total energy, $E[\varphi_{\text{n}}]=E[\varphi_{\text{c}}]$. However, recall that the angular numbers must satisfy the conditions below:
@@
From the fact that $\Lambda_{\text{n}}=\Lambda_{\text{c}}$ it is clear that $\ell_{\text{n}}<\ell_{\text{c}}$, so in the charged case $\lvert m_{\text{c}}\rvert$ can take larger values than $\lvert m_{\text{n}}\rvert$. Thus, the charged field can carry more angular momentum, as follows from Eq. (75). In particular, focusing on the field states that possess a maximum amount of angular momentum, $\ell_{\text{n,c}}=m_{\text{n,c}}$, we conclude that
@@
To explain this difference, it is useful to illustrate how $\varphi_{\text{n,c}}$ distribute in space. For concreteness, we shall restrict ourselves to fundamental modes with zero KK momentum and take the same background parameters and field mass of Fig. 6 ($\eta=1.75$, $\tilde{\mu}=7$). For $\varphi_{\text{n}}$, we shall take maximum angular momentum states with $\ell_{\text{n}}=m_{\text{n}}=6,7,8$, which according to the mode classification given in Section III.2 correspond to TS, mixed and cloud modes, respectively. To guarantee that $\Lambda_{\text{n}}=\Lambda_{\text{c}}$, in the charged case we also take maximum angular momentum states with $\ell_{\text{c}}=m_{\text{c}}=N_{\text{c}}/2$ and $N_{\text{c}}=84,112,144$. We emphasize that the condition $m_{\text{c}}=N_{\text{c}}/2$ corresponds to a north monopole mode, that is, a mode that is localized along the north component of the $z$-axis and carries angular momentum in the $z$-direction Pereñiguez et al. (2024) (see also Appendix A). In all cases, we normalize the wavefunctions by setting the total energy to $E[\varphi_{\text{n,c}}]/r_{B}=10^{-3}$.